OpenStudy (anonymous):

what is 8 over 3 times 36

OpenStudy (anonymous):

8/3*36 = 8/108 = 4/54 = 2/27

OpenStudy (anonymous):

I have a question for clarification.. is it (8/3) * 36 or 8/(3*36)

OpenStudy (anonymous):

I think it might be (8/3) * 36